# Haptic Perception via the Dynamics of Flexible Body Inspired by an Ostrich’s Neck

### v1

@@ Abstract
In biological systems, haptic perception is achieved through both flexible skin and flexible body. In fully soft robots, the fragility of their bodies and the time delays in sensory processing pose significant challenges. The musculoskeletal system possesses both the deformability inherent in soft materials and the durability of rigid-body robots. Additionally, by outsourcing part of the intelligent information processing to the morphology of the musculoskeletal system, applications for dynamic tasks are expected. This study focuses on the pecking movements of birds, which achieve precise haptic perception through the musculoskeletal system of their flexible neck. Physical reservoir computing is applied to flexible structures inspired by an ostrich neck to analyze the relationship between haptic perception and physical characteristics. Combined experiments using both an actual robot and simulations demonstrate that, under appropriate body viscoelasticity, the flexible structure can distinguish objects of varying softness and memorize this information as behaviors. Drawing on these findings and anatomical insights from the ostrich neck, a haptic sensing system is proposed that possesses separability and this behavioral memory in flexible structures, enabling rapid learning and real-time inference. The results demonstrate that through the dynamics of flexible structures, diverse functions can emerge beyond their original design as manipulators.
@@
Physical Reservoir Computing, Haptic Perception, Bio-inspired Robotics
@@ 1 Introduction
The musculoskeletal system found in many organisms can achieve a balance between the continuous deformability inherent in soft robots and the robustness, high output and high precision characteristics of rigid robots[24]. In the musculoskeletal system, constraints due to morphological structures such as joint structure and muscle-tendon arrangements can facilitate some intelligent behaviors. For example, the viscoelasticity and anisotropic compliance inherent in the musculoskeletal system were utilized to perform complex and dynamic tasks, such as identifying objects and opening doors, using simple control inputs[25]. In this case, some of the calculations necessary for control are outsourced to the physical body itself[26]. This framework is referred to as morphological computation[27], and it has been shown that utilizing it enables intelligent behaviors such as gait, with fewer computational resources[28, 29]. We consider the use of the musculoskeletal system and its dynamics to be one of the most promising approaches to take. Therefore, we focus on the pecking behavior of birds, which achieves precise tactile sensing using the flexible musculoskeletal system of the neck[30]. As shown in Figure 1(a), we dissected an ostrich’s neck and identified the muscle-tendon arrangement for its typical pecking movement, and developed a flexible robot capable of performing pecking movements similar to those of an ostrich in our previous researches[31, 32]. Furthermore, we confirmed that the robot exhibited a variety of dynamics (i.e., time series of joint angles) in response to different external forces, due to its passivity and multi-degree-of-freedom characteristics[33]. However, the behaviors were highly sensitive to physical characteristics. In morphological computation, it has been reported that input and body parameters have a significant impact on the resulting behaviors and tactile sensation[34, 16, 35]. A comprehensive understanding of the relationship between input parameters, physical characteristics, and haptic perception is required.
@@
  • Comprehensive simulator experiments revealed that the dynamics of a flexible body possess the ability to amplify differences in external forces (referred to as separability) and to retain these differences as behavior (referred to as haptic memory).
@@
  • We found that morphological features, such as elastic elements for gravity compensation and the heterogeneity of body viscoelasticity, enhance both separability and haptic memory.
  • The proposed system uses physical reservoir computing to seamlessly perform data acquisition and learning (which takes about 25 seconds in our experiments). In addition, it can infer the softness of collided objects in real time.
  • The body dynamics of biologically inspired flexible structures, originally developed as manipulators, have been shown to have functions similar to those of sequential circuits and can reduce the computational load.
@@ 2.1 Physical Reservoir Computing
Morphological computation views the complex dynamics of flexible structures not only as a control problem but also as a computational resource[27]. This computational capability can be formulated as the ability to emulate the mapping from input to output when a static and linear readout is added to a morphology such as a mass-spring model[36]. This method is commonly referred to as physical reservoir computing[37]. In this approach, physical systems, such as the body of a soft robot, are regarded a recurrent neural network with fixed weights. By training the readout layer, it is possible to emulate desired functions such as sensing and control[38, 39]. In physical reservoir computing, two inherent capabilities of physical systems are leveraged: the ability to non-linearly map low-dimensional inputs to a high-dimensional space, similar to kernels in machine learning, and the ability to temporarily retain the influence of recent input sequences[40]. Previous studies have thoroughly analyzed the positional distribution of computational capabilities that originate from the non-linearity and memory inherent in the body dynamics of soft robots[41]. In the context of haptic perception in a flexible body, methods have also been developed to quantify and analyze the spatio-temporal propagation of information related to contact in a flexible body[42]. These studies show that the framework of physical reservoir computing and its analysis is useful to quantify the contribution of body parameters, morphology, and behavior to perception[Luís2019].
@@ 2.3 Behaviors that Enable Haptic Perception
Figure 3(a) shows a heat map that illustrates the relationship between body stiffness $S$ and input speed $A/T$ at the maximum value of learning performance. Figure 3(b) shows the behaviors and the time series of the reaction force at the maximum learning performance for each input speed. The supporting information provides the learning performance for all patterns (12 patterns for input parameters and 225 patterns for body parameters), along with the corresponding behavior and time series data. Figure 3 (a) shows that learning performance is low for inputs that are either too rapid (240 rad/s) or too slow (15 rad/s). As shown in Figure 3 (b), at 240 rad/s, the input exceeds the viscoelastic capacity of the body, disrupting stable pecking. Additionally, at 15 rad/s, the generated reaction force is insufficient to obtain meaningful information about the pecked object. These observations indicate that physical constraints, such as the inability to maintain stable pecking, and informational constraints, such as insufficient reaction force to extract information from the environment, both play crucial roles in enabling learning. In these examples, the conditions that reduce learning performance are clear. Interestingly, even under seemingly favorable conditions such as 120 rad/s—with stable pecking and stronger reaction force than at 15 rad/s—the learning performance remains lower, suggesting more subtle underlying physical and informational constraints. Even in cases where the possibility of learning is not clear from the behavior, physical and informational conditions are likely to exist. Focusing on conditions with high learning performance (45-90 rad/s), as the input speed increases, the value of body stiffness that corresponds to the maximum value of the heat map decreases. In other words, there are strategies that involve driving a flexible body at high speed or a rigid body at low speed. The behavior under each condition is shown in the supporting information video 2.
@@ 2.5 Separability
In the following, these findings are validated through actual robot experiments. First, we compare the change in accuracy with respect to the input amplitude $A$. A larger input amplitude corresponds to a rapid input speed. Figure 5(d) shows the accuracy curve. The x-axis represents the time over one pecking cycle and the y-axis represents the accuracy. The black line in the accuracy curve of Figure 5(d) represents the accuracy under conditions with a relatively rapid input speed (e.g., $A/T$ = 80 deg/s), while the colored lines represent conditions with a slow input speed (e.g., $A/T$ = 20 deg/s). The black line shows that, under the rapid input condition, the learning performance remains low, even when the object is periodically colliding with the system. The center part of Figure 5(d) shows the ESP index and the maximum values of the silhouette coefficient for each case. When the input speed is high, the ESP index increases, indicating that the ESP does not hold. In other words, by focusing on the reproducibility of the joint angle time series, we can determine the input parameters that enable learning. Next, we examine accuracy changes by adjusting the viscosity with a friction brake, while keeping the input parameters fixed under conditions where the ESP holds. As shown in the accuracy curve, learning performance is low when the body viscosity is either too high or too low. Additionally, as shown in the center part of Figure 5(d), the learning performance is highest under conditions where the silhouette coefficient are maximized. This means that by focusing on the separation and deviation of rebound forces, we can estimate the learning performance. The right part of Figure 5(d) shows the time series of the reactive force and the silhouette coefficient for proper body viscoelasticity. This figure illustrates that, in actual robot experiments, the separation of the trajectory occurs during the bounce that follows the impact force. By focusing on the reproducibility of body dynamics (Echo State Property) and the separability and deviation of the reactive force time series, we can set appropriate input and body parameters before training and evaluation. The data from these actual robot experiments are shown in the supporting information. Furthermore, we measured the actual pecking reaction forces of the ostrich and confirmed the presence of reaction force time series with multiple peaks, rather than unimodal ones (see the supporting information).
@@ 3 Conclusion
Next, we demonstrated that under conditions of rapid input and flexible body parameters, information from the previous cycle of periodic movements can be retained as behavior. We quantified this memory using physical reservoir computing and showed that incorporating the heterogeneity of body viscoelasticity, inspired by the arrangement of muscles and ligaments in an ostrich’s neck, enhances this memory. The evaluation of computational capabilities of biological anatomical structures using reservoir computing is gaining attention as a promising approach for mutual inspiration between biology and computational science[Luís2019, 61]. The actual ostrich neck consists of about 200 muscles and a complex system of force transmission through tissues with varying material properties[48, 62], such as fascia and tendons, suggesting the potential for even more diverse computational capabilities. Our results represent an initial step in exploring this potential. Based on these findings, we propose a haptic perception system that enables rapid learning and real-time inference by utilizing musculoskeletal dynamics. This task is analogous to how bipedal animals recognize ground conditions during running and hopping[60, 63] . In such cases, animals must rapidly adapt to changes in surface properties under the constraints of sensorimotor delays (e.g., nerve transmission delays and muscle electromechanical coupling) to prevent falling[64]. The information from the previous cycle, retained within the body, is considered useful for generating predictive feedforward commands that enable adaptive behavior. In this study, the input and body parameters for haptic perception were predefined manually. Therefore, in future work, we propose to build a closed-loop system and develop mechanisms and algorithms for parameter adjustment. Recently, Augmented Direct Feedback Alignment, which replaces state updates based on the gradient of nonlinear mappings with random matrices, has been proposed[65]. We are considering using this algorithm to adjust the body parameters.
@@ Acknowledgements
This work is supported by JSPS KAKENHI Grant Numbers 24KJ0599 and 21KK0182 and 23K18472 and by JST CREST Grant Number JPMJCR2014. K. I. was supported by JSPS KAKENHI Grant Number 22K21295. The authors thank Dr. Megu Gunji and Dr. Mitsuhito Ando for providing the data and photographs used in this article.

### v2

@@ Abstract
In biological systems, both skin sensitivity and body flexibility play crucial roles in haptic perception. Fully soft robots often suffer from structural fragility and delayed sensory processing, limiting their practical functionality. The musculoskeletal system combines the adaptability of soft materials with the durability of rigid-body robots. It also leverages morphological computation, where the morphological structures contribute to information processing, for dynamic and adaptive behaviors. This study focuses on the pecking behaviors of birds, which enables precise haptic perception through the musculoskeletal system of their flexible neck. Physical reservoir computing is applied to flexible structures inspired by an ostrich neck to analyze the relationship between haptic perception and physical characteristics. Experiments with both a physical robot and simulations reveal that, with appropriate viscoelasticity, the flexible structure can discriminate object softness and retain that information through behavior. Drawing on these findings and anatomical insights from the ostrich neck, a haptic perception system is proposed that exhibits both separability and behavioral memory in flexible structures, enabling rapid learning and real-time inference. The results demonstrate that through the dynamics of flexible structures, diverse functions can emerge beyond their original design as manipulators.
@@
Biologically Inspired Robotics, Morphological Computation, Soft Robotics, Physical Reservoir Computing, Embodied Intelligence
@@ 1 Introduction
The musculoskeletal system found in many organisms can achieve a balance between the continuous deformability inherent in soft robots and the robustness, high output and high precision characteristics of rigid robots[24]. In the musculoskeletal system, constraints due to morphological structures such as joint structure and muscle-tendon arrangements can facilitate some intelligent behaviors. For example, the viscoelasticity and anisotropic compliance inherent in the musculoskeletal system were utilized to perform complex and dynamic tasks, such as identifying objects and opening doors, using simple control inputs[25]. In this case, some of the calculations necessary for control are outsourced to the physical body itself[26]. This framework is referred to as morphological computation[27], and it has been shown that utilizing it enables intelligent behaviors such as gait, with fewer computational resources[28, 29]. We consider the use of the musculoskeletal system and its dynamics to be one of the most promising approaches to take. Therefore, we focus on the pecking behavior of birds, which achieves precise tactile sensing using the flexible musculoskeletal system of the neck[30]. As shown in Figure 1(a), we dissected an ostrich’s neck and identified the muscle-tendon arrangement for its typical pecking behavior, and developed a flexible robot capable of performing pecking behaviors similar to those of an ostrich in our previous researches[31, 32]. (Supporting information video 1 shows the behavior of RobOstrich in detail.) Furthermore, we confirmed that the robot exhibited a variety of dynamics (i.e., time series of joint angles) in response to different external forces, due to its passivity and multi-degree-of-freedom characteristics[33]. However, the behaviors were highly sensitive to physical characteristics. In morphological computation, it has been reported that input and body parameters have a significant impact on the resulting behaviors and tactile sensation[34, 16, 35]. A comprehensive understanding of the relationship between input parameters, physical characteristics, and haptic perception is required.
@@
  • Experiments using both a physical robot and simulations demonstrate that the dynamics of a flexible body can amplify subtle differences in external forces—termed separability—and preserve them as behavioral patterns—termed haptic memory.
@@
  • We evaluated the effect of heterogeneity in body viscoelasticity, resulting from morphological structures such as ligament and muscle arrangements, on both separability and haptic memory.
  • The system leverages physical reservoir computing for seamless data acquisition and learning, completing the training in approximately 25 seconds. It further enables real-time inference of the softness of collided objects.
  • We demonstrate that this haptic memory allows for state-dependent outputs, like a sequential circuit, and can reduce the computational load in sensor data processing.
These findings indicate that leveraging the inherent dynamics of a flexible body can lead to emergent functionalities beyond the original design intentions as a manipulator, highlighting their potential in biologically inspired soft robotic systems.
@@ 2.1 Physical Reservoir Computing
Morphological computation views the complex dynamics of flexible structures not only as a control problem but also as a computational resource[27]. This computational capability can be formulated as the ability to emulate the mapping from input to output when a static and linear readout is added to a morphology such as a mass-spring model[36]. This method is commonly referred to as physical reservoir computing[37]. In this approach, physical systems, such as the body of a soft robot, are regarded a recurrent neural network with fixed weights. By training the readout layer, it is possible to emulate desired functions such as sensing and control[38, 39]. In physical reservoir computing, two inherent capabilities of physical systems are leveraged: the ability to non-linearly map low-dimensional inputs to a high-dimensional space, similar to kernels in machine learning, and the ability to temporarily retain the influence of recent input sequences[40]. Previous studies have thoroughly analyzed the positional distribution of computational capabilities that originate from the non-linearity and memory inherent in the body dynamics of soft robots[41]. In the context of haptic perception in a flexible body, methods have also been developed to quantify and analyze the spatio-temporal propagation of information related to contact in a flexible body[42]. These studies show that the framework of physical reservoir computing and its analysis is useful to quantify the contribution of body parameters, morphology, and behavior to perception[43].
@@ 2.3 Behaviors that Enable Haptic Perception
Figure 3(a) shows a heat map that illustrates the relationship between body stiffness $S$ and input speed $A/T$ at the maximum value of learning performance. Figure 3(b) shows the behaviors and the time series of the reaction force at the maximum learning performance for each input speed. The supporting information provides the learning performance for all patterns (12 patterns for input parameters and 225 patterns for body parameters), along with the corresponding behavior and time series data. Figure 3 (a) shows that learning performance is low for inputs that are either too rapid (240 rad/s) or too slow (15 rad/s). As shown in Figure 3 (b), at 240 rad/s, the input exceeds the viscoelastic capacity of the body, disrupting stable pecking. Additionally, at 15 rad/s, the generated reaction force is insufficient to obtain meaningful information about the pecked object. These observations indicate that physical conditions, such as the inability to maintain stable pecking, and informational conditions, such as insufficient reaction force to extract information from the environment, both play crucial roles in enabling learning. In these examples, the conditions that reduce learning performance are clear. Notably, even under seemingly favorable conditions such as 120 rad/s—with stable pecking and stronger reaction force than at 15 rad/s—the learning performance remains lower, suggesting more subtle underlying physical and informational conditions. Even in cases where the possibility of learning is not clear from the behavior, physical and informational conditions are likely to exist. Focusing on conditions with high learning performance (45-90 rad/s), as the input speed increases, the value of body stiffness that corresponds to the maximum value of the heat map decreases. In other words, there are strategies that involve driving a flexible body at high speed or a rigid body at low speed. The diversity of these strategies is thought to contribute to adaptability, allowing the behavior to be modulated according to different environmental conditions or task demands[53]. The behavior under each condition is shown in the supporting information video 2.
@@ 2.5 Separability
In the following, these findings are validated through actual robot experiments. First, we compare the change in accuracy with respect to the input amplitude $A$. A larger input amplitude corresponds to a rapid input speed. Figure 5(d) shows the accuracy curve. The x-axis represents the time over one pecking cycle and the y-axis represents the accuracy. The black line in the accuracy curve of Figure 5(d) represents the accuracy under conditions with a relatively rapid input speed (e.g., $A/T$ = 80 deg/s), while the colored lines represent conditions with a slow input speed (e.g., $A/T$ = 20 deg/s). The black line shows that, under the rapid input condition, the learning performance remains low, even when the object is periodically colliding with the system. The center part of Figure 5(d) shows the ESP index and the maximum values of the silhouette coefficient for each case. When the input speed is high, the ESP index increases, indicating that the ESP does not hold. In other words, by focusing on the reproducibility of the joint angle time series, we can determine the input parameters that enable learning. Next, we examine accuracy changes by adjusting the viscosity with a friction brake, while keeping the input parameters fixed under conditions where the ESP holds. As shown in the accuracy curve, learning performance is low when the body viscosity is either too high or too low. Additionally, as shown in the center part of Figure 5(d), the learning performance is highest under conditions where the silhouette coefficient are maximized. This means that by focusing on the separation and deviation of rebound forces, we can estimate the learning performance. The right part of Figure 5(d) shows the time series of the reactive force and the silhouette coefficient for proper body viscoelasticity. This figure illustrates that, in actual robot experiments, the separation of the trajectory occurs during the bounce that follows the impact force. By focusing on the reproducibility of body dynamics (Echo State Property) and the separability and deviation of the reactive force time series, we can set appropriate input and body parameters before training and evaluation. The data from these actual robot experiments are shown in the supporting information.
@@ 3 Conclusion
Next, we demonstrated that under conditions of rapid input and flexible body parameters, information from the previous cycle of periodic movements can be retained as behavior. We quantified this memory using physical reservoir computing and showed that incorporating the heterogeneity of body viscoelasticity, inspired by the arrangement of muscles and ligaments in an ostrich’s neck, enhances this memory. The evaluation of computational capabilities of biological anatomical structures using reservoir computing is gaining attention as a promising approach for mutual inspiration between biology and computational science[43, 62]. The actual ostrich neck consists of about 200 muscles and a complex system of force transmission through tissues with varying material properties[48, 63], such as fascia and tendons, suggesting the potential for even more diverse computational capabilities. Our results represent an initial step in exploring this potential. Based on these findings, we propose a haptic perception system that enables rapid learning and real-time inference by utilizing musculoskeletal dynamics. This task is analogous to how bipedal animals recognize ground conditions during running and hopping[61, 64] . In such cases, animals must rapidly adapt to changes in surface properties under the constraints of sensorimotor delays (e.g., nerve transmission delays and muscle electromechanical coupling) to prevent falling[65]. The information from the previous cycle, retained within the body, is considered useful for generating predictive feedforward commands that enable adaptive behavior. In this study, the input and body parameters for haptic perception were predefined manually. Therefore, in future work, we propose to build a closed-loop system and develop mechanisms and algorithms for parameter adjustment. Recently, Augmented Direct Feedback Alignment, which replaces state updates based on the gradient of nonlinear mappings with random matrices, has been proposed[66]. We are considering using this algorithm to adjust the body parameters.
@@ Acknowledgements
This work is supported by JSPS KAKENHI Grant Numbers 24KJ0599 and 21KK0182 and 23K18472 and by JST CREST Grant Number JPMJCR2014. K. I. was supported by JSPS KAKENHI Grant Number 22K21295. The authors thank Dr. Megu Gunji for providing photographs used in this article.
Ethics Approval
This paper includes anatomical illustrations of an ostrich. The specimen used was obtained from an animal processed for food purposes, not specifically for research, and therefore its use does not involve any ethical issues related to animal experimentation.